# Block triangular preconditioning for inverse source problems in time-space fractional diffusion equations

### v1

@@ 6.4 Accuracy of the reconstruction
In the first row, corresponding to a low noise level $\varepsilon=0.001$, the reconstruction achieves high accuracy when $\lambda=10^{-4}$. A smaller value, $\lambda=10^{-5}$, results in overfitting to noise, introducing oscillations. On the other hand, with a larger value $\lambda=10^{-3}$ we begin to see an oversmoothing effect. As the noise level increases to $\varepsilon=0.01$ (second row), the reconstruction becomes more challenging. Nonetheless, $\lambda=10^{-3}$ provides a reasonable balance between fidelity and regularity. With the other two values we obtain similar deficiencies as before: $\lambda=10^{-4}$ introduces oscillations, while $\lambda=10^{-2}$ causes oversmoothing. At the highest noise level, $\varepsilon=0.1$ (third row), accurate reconstruction becomes significantly more difficult. Among the tested values, $\lambda=10^{-2}$ yields the most reliable approximation, whereas both $\lambda=10^{-3}$ and $\lambda=10^{-1}$ produce reconstructions that deviate substantially from the exact source. Similar patterns are observed in Figures 3 and 4, which depict the results for the fractional orders $\beta=0.5$, $\omega=1.5$ and $\beta=0.9$, $\omega=1.1$, respectively.
@@
Figure 2: Comparison between the exact source function $f(x)$, in blue, and its reconstruction, in red, for different choices of the regularization parameter $\lambda$ and noise level $\varepsilon$. The problem sizes are $n=S=200$ and the fractional orders are $\beta=0.1$, $\omega=1.9$.
@@
Figure 3: Comparison between the exact source function $f(x)$, in blue, and its reconstruction, in red, for different choices of the regularization parameter $\lambda$ and noise level $\varepsilon$. The problem sizes are $n=S=200$ and the fractional orders are $\beta=0.5$, $\omega=1.5$.
@@
Figure 4: Comparison between the exact source function $f(x)$, in blue, and its reconstruction, in red, for different choices of the regularization parameter $\lambda$ and noise level $\varepsilon$. The problem sizes are $n=S=200$ and the fractional orders are $\beta=0.9$, $\omega=1.1$.
@@
Table 3: 2-norm relative error between the exact source function $f(x)$ and its reconstruction, for different choices of the regularization parameter $\lambda$ and noise level $\varepsilon$. The fractional orders are $\beta=0.1$, $\omega=1.9$.
$\varepsilon=0.001$ $\varepsilon=0.01$ $\varepsilon=0.1$
$\lambda$ $\lambda$ $\lambda$
$n$ $S$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-3}$ $10^{-2}$ $10^{-1}$
$2^{4}$ $2^{4}$ 0.0165 0.0133 0.0506 0.1328 0.0721 0.2571 0.3573 0.2108 0.6351
$2^{4}$ $2^{5}$ 0.0234 0.0167 0.0540 0.1083 0.0614 0.2358 0.3515 0.1995 0.6208
$2^{4}$ $2^{6}$ 0.0288 0.0160 0.0515 0.1264 0.0711 0.2504 0.3176 0.2071 0.6390
$2^{4}$ $2^{7}$ 0.0213 0.0153 0.0525 0.1026 0.0747 0.2448 0.3523 0.2139 0.6126
$2^{4}$ $2^{8}$ 0.0328 0.0140 0.0515 0.1353 0.0634 0.2424 0.3519 0.2366 0.6545
$2^{5}$ $2^{4}$ 0.0548 0.0170 0.0527 0.1566 0.0848 0.2498 0.3907 0.2321 0.6119
$2^{5}$ $2^{5}$ 0.0438 0.0173 0.0526 0.1241 0.0594 0.2400 0.4186 0.3003 0.6227
$2^{5}$ $2^{6}$ 0.0562 0.0185 0.0516 0.1197 0.0541 0.2497 0.2448 0.1855 0.6347
$2^{5}$ $2^{7}$ 0.0454 0.0155 0.0541 0.1458 0.0612 0.2506 0.4665 0.1576 0.6347
$2^{5}$ $2^{8}$ 0.0460 0.0156 0.0532 0.1420 0.0473 0.2455 0.3507 0.3125 0.5993
$2^{6}$ $2^{4}$ 0.0514 0.0177 0.0500 0.1049 0.0697 0.2396 0.3966 0.2701 0.5992
$2^{6}$ $2^{5}$ 0.0592 0.0163 0.0524 0.1489 0.0441 0.2394 0.3684 0.2861 0.6170
$2^{6}$ $2^{6}$ 0.0579 0.0194 0.0510 0.1628 0.0552 0.2393 0.5085 0.3732 0.6115
$2^{6}$ $2^{7}$ 0.0658 0.0154 0.0513 0.1308 0.0452 0.2531 0.3390 0.2057 0.6130
$2^{6}$ $2^{8}$ 0.0546 0.0181 0.0525 0.1976 0.0637 0.2349 0.3079 0.1814 0.6516
$2^{7}$ $2^{4}$ 0.0680 0.0138 0.0517 0.1734 0.0817 0.2451 0.3232 0.3083 0.6233
$2^{7}$ $2^{5}$ 0.0618 0.0175 0.0511 0.1422 0.0692 0.2530 0.3489 0.4290 0.6372
$2^{7}$ $2^{6}$ 0.0591 0.0166 0.0543 0.1136 0.0779 0.2499 0.3620 0.2395 0.5929
$2^{7}$ $2^{7}$ 0.0662 0.0199 0.0510 0.1517 0.0625 0.2400 0.3291 0.3030 0.6373
$2^{7}$ $2^{8}$ 0.0570 0.0139 0.0532 0.1406 0.0561 0.2517 0.3791 0.3510 0.6376
$2^{8}$ $2^{4}$ 0.0665 0.0221 0.0528 0.1609 0.0635 0.2409 0.4625 0.2889 0.6152
$2^{8}$ $2^{5}$ 0.0719 0.0163 0.0537 0.1643 0.0782 0.2412 0.4098 0.3255 0.5931
$2^{8}$ $2^{6}$ 0.0621 0.0177 0.0507 0.1789 0.0762 0.2363 0.3899 0.2234 0.6276
$2^{8}$ $2^{7}$ 0.0659 0.0123 0.0541 0.1430 0.0774 0.2408 0.4459 0.3415 0.6030
@@
Table 4: 2-norm relative error between the exact source function $f(x)$ and its reconstruction, for different choices of the regularization parameter $\lambda$ and noise level $\varepsilon$. The fractional orders are $\beta=0.5$, $\omega=1.5$.
$\varepsilon=0.001$ $\varepsilon=0.01$ $\varepsilon=0.1$
$\lambda$ $\lambda$ $\lambda$
$n$ $S$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-3}$ $10^{-2}$ $10^{-1}$
$2^{4}$ $2^{4}$ 0.0130 0.0090 0.0397 0.1050 0.0460 0.2186 0.3252 0.2125 0.6136
$2^{4}$ $2^{5}$ 0.0082 0.0096 0.0396 0.0766 0.0440 0.2289 0.4063 0.1563 0.6605
$2^{4}$ $2^{6}$ 0.0164 0.0109 0.0396 0.0704 0.0485 0.2294 0.4200 0.3256 0.6539
$2^{4}$ $2^{7}$ 0.0176 0.0077 0.0403 0.1118 0.0480 0.2316 0.3140 0.2256 0.5865
$2^{4}$ $2^{8}$ 0.0108 0.0096 0.0400 0.0893 0.0650 0.2116 0.3861 0.2221 0.6202
$2^{5}$ $2^{4}$ 0.0322 0.0131 0.0381 0.1701 0.0585 0.2191 0.4615 0.2271 0.6263
$2^{5}$ $2^{5}$ 0.0389 0.0187 0.0376 0.1390 0.0587 0.2274 0.4847 0.1752 0.5671
$2^{5}$ $2^{6}$ 0.0308 0.0166 0.0365 0.1512 0.0575 0.2274 0.4104 0.2353 0.6506
$2^{5}$ $2^{7}$ 0.0392 0.0164 0.0389 0.1303 0.0732 0.2179 0.4401 0.1031 0.6224
$2^{5}$ $2^{8}$ 0.0347 0.0145 0.0366 0.1889 0.0619 0.2243 0.4104 0.2872 0.6204
$2^{6}$ $2^{4}$ 0.0678 0.0194 0.0397 0.1914 0.0587 0.2144 0.4517 0.1754 0.6314
$2^{6}$ $2^{5}$ 0.0686 0.0199 0.0407 0.1819 0.0570 0.2312 0.5112 0.2453 0.5970
$2^{6}$ $2^{6}$ 0.0639 0.0197 0.0383 0.2132 0.0648 0.2230 0.5637 0.2786 0.6228
$2^{6}$ $2^{7}$ 0.0610 0.0191 0.0395 0.1708 0.0505 0.2234 0.4392 0.3005 0.6163
$2^{6}$ $2^{8}$ 0.0632 0.0181 0.0376 0.1844 0.0614 0.2280 0.5383 0.3175 0.6296
$2^{7}$ $2^{4}$ 0.0888 0.0184 0.0375 0.2020 0.0490 0.2288 0.5170 0.3455 0.6303
$2^{7}$ $2^{5}$ 0.0805 0.0238 0.0392 0.1875 0.0574 0.2217 0.4128 0.3175 0.6304
$2^{7}$ $2^{6}$ 0.0774 0.0212 0.0402 0.1873 0.0556 0.2285 0.4635 0.1273 0.6373
$2^{7}$ $2^{7}$ 0.0902 0.0210 0.0377 0.1980 0.0523 0.2213 0.5139 0.2577 0.6487
$2^{7}$ $2^{8}$ 0.0792 0.0225 0.0395 0.1848 0.0674 0.2309 0.4703 0.2019 0.6404
$2^{8}$ $2^{4}$ 0.0926 0.0214 0.0395 0.2351 0.0696 0.2211 0.6011 0.1947 0.6014
$2^{8}$ $2^{5}$ 0.0985 0.0237 0.0400 0.2267 0.0923 0.2131 0.4037 0.2442 0.6338
$2^{8}$ $2^{6}$ 0.0926 0.0216 0.0392 0.1813 0.0541 0.2222 0.5407 0.2804 0.6536
$2^{8}$ $2^{7}$ 0.0868 0.0203 0.0394 0.2278 0.0611 0.2188 0.5782 0.1801 0.6230
@@
Table 5: 2-norm relative error between the exact source function $f(x)$ and its reconstruction, for different choices of the regularization parameter $\lambda$ and noise level $\varepsilon$. The fractional orders are $\beta=0.9$, $\omega=1.1$.
$\varepsilon=0.001$ $\varepsilon=0.01$ $\varepsilon=0.1$
$\lambda$ $\lambda$ $\lambda$
$n$ $S$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-4}$ $10^{-3}$ $10^{-2}$ $10^{-3}$ $10^{-2}$ $10^{-1}$
$2^{4}$ $2^{4}$ 0.0064 0.0059 0.0384 0.0548 0.0427 0.2396 0.3255 0.1787 0.6776
$2^{4}$ $2^{5}$ 0.0069 0.0056 0.0385 0.0426 0.0576 0.2357 0.3794 0.1927 0.6632
$2^{4}$ $2^{6}$ 0.0038 0.0068 0.0390 0.0302 0.0521 0.2335 0.2868 0.2326 0.6365
$2^{4}$ $2^{7}$ 0.0060 0.0040 0.0385 0.0627 0.0320 0.2369 0.2878 0.3304 0.6194
$2^{4}$ $2^{8}$ 0.0063 0.0074 0.0386 0.0559 0.0520 0.2345 0.3208 0.3305 0.6612
$2^{5}$ $2^{4}$ 0.0114 0.0109 0.0378 0.0811 0.0465 0.2300 0.3843 0.2132 0.6175
$2^{5}$ $2^{5}$ 0.0124 0.0128 0.0382 0.1242 0.0592 0.2402 0.4434 0.2255 0.6383
$2^{5}$ $2^{6}$ 0.0155 0.0102 0.0376 0.1071 0.0627 0.2190 0.3545 0.3347 0.6544
$2^{5}$ $2^{7}$ 0.0148 0.0080 0.0383 0.0862 0.0732 0.2323 0.4037 0.2575 0.6527
$2^{5}$ $2^{8}$ 0.0097 0.0134 0.0375 0.1110 0.0632 0.2299 0.3936 0.1676 0.6301
$2^{6}$ $2^{4}$ 0.0345 0.0194 0.0394 0.1766 0.0659 0.2245 0.4922 0.3359 0.6176
$2^{6}$ $2^{5}$ 0.0255 0.0170 0.0380 0.1610 0.0582 0.2441 0.6122 0.2556 0.6742
$2^{6}$ $2^{6}$ 0.0311 0.0155 0.0387 0.1792 0.0960 0.2397 0.4798 0.2300 0.6336
$2^{6}$ $2^{7}$ 0.0284 0.0161 0.0384 0.1672 0.0634 0.2295 0.5026 0.1870 0.6377
$2^{6}$ $2^{8}$ 0.0337 0.0190 0.0393 0.1776 0.0635 0.2221 0.5682 0.2073 0.6382
$2^{7}$ $2^{4}$ 0.0718 0.0251 0.0392 0.2170 0.0756 0.2201 0.6129 0.3203 0.6746
$2^{7}$ $2^{5}$ 0.0632 0.0253 0.0396 0.2275 0.0752 0.2295 0.6186 0.3296 0.6241
$2^{7}$ $2^{6}$ 0.0696 0.0242 0.0372 0.2487 0.0624 0.2402 0.5921 0.3867 0.6566
$2^{7}$ $2^{7}$ 0.0618 0.0264 0.0384 0.2250 0.0867 0.2353 0.5685 0.2039 0.6543
$2^{7}$ $2^{8}$ 0.0698 0.0247 0.0403 0.2355 0.0670 0.2292 0.5926 0.1622 0.6517
$2^{8}$ $2^{4}$ 0.1107 0.0314 0.0385 0.3278 0.0766 0.2131 0.6623 0.2643 0.6353
$2^{8}$ $2^{5}$ 0.1053 0.0312 0.0405 0.2877 0.0844 0.2360 0.7304 0.3000 0.6684
$2^{8}$ $2^{6}$ 0.1007 0.0295 0.0398 0.2935 0.0742 0.2261 0.7023 0.2208 0.5910
$2^{8}$ $2^{7}$ 0.1036 0.0295 0.0394 0.2804 0.0840 0.2330 0.7155 0.2104 0.6462

### v2

@@ 6.1.3 Accuracy of the reconstruction
Figure 2 shows the function $f(x)$ and its reconstruction under varying noise levels $\varepsilon$ and choices of regularization parameter $\lambda$, with the fractional orders fixed at $\beta=0.1$, $\omega=1.9$ and sizes $n=2^{8}$, $S=2^{6}$. In the first row, corresponding to a low noise level $\varepsilon=0.001$, the reconstruction achieves high accuracy when $\lambda=10^{-4}$. A smaller value, $\lambda=10^{-5}$, results in overfitting to noise, introducing oscillations. On the other hand, with a larger value $\lambda=10^{-3}$ we begin to see an oversmoothing effect. As the noise level increases to $\varepsilon=0.01$ (second row), the reconstruction becomes more challenging. Nonetheless, $\lambda=10^{-3}$ provides a reasonable balance between fidelity and regularity. With the other two values we obtain similar deficiencies as before: $\lambda=10^{-4}$ introduces oscillations, while $\lambda=10^{-2}$ causes oversmoothing.
Similar patterns are observed in Figures 3 and 4, which depict the results for the fractional orders $\beta=0.5$, $\omega=1.5$ and $\beta=0.9$, $\omega=1.1$, respectively.
@@
Figure 2: 1D case - Comparison between the exact source function $f(x)$, in blue, and its reconstruction, in red, for different choices of the regularization parameter $\lambda$ and noise level $\varepsilon$. The problem sizes are $n=2^{8}$, $S=2^{6}$ and the fractional orders are $\beta=0.1$, $\omega=1.9$.
@@
Figure 3: 1D case - Comparison between the exact source function $f(x)$, in blue, and its reconstruction, in red, for different choices of the regularization parameter $\lambda$ and noise level $\varepsilon$. The problem sizes are $n=2^{8}$, $S=2^{6}$ and the fractional orders are $\beta=0.5$, $\omega=1.5$.
@@
Figure 4: 1D case - Comparison between the exact source function $f(x)$, in blue, and its reconstruction, in red, for different choices of the regularization parameter $\lambda$ and noise level $\varepsilon$. The problem sizes are $n=2^{8}$, $S=2^{6}$ and the fractional orders are $\beta=0.9$, $\omega=1.1$.
@@
Table 3: 1D case - 2-norm relative error between the exact source function $f(x)$ and its reconstruction, for different choices of the regularization parameter $\lambda$ and noise level $\varepsilon$. The fractional orders are $\beta=0.1$, $\omega=1.9$.
$\varepsilon=0.001$ $\varepsilon=0.01$
$\lambda$ $\lambda$
$n$ $S$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-4}$ $10^{-3}$ $10^{-2}$
$2^{4}$ $2^{4}$ 0.02 0.01 0.05 0.11 0.07 0.26
$2^{4}$ $2^{5}$ 0.03 0.01 0.05 0.10 0.07 0.25
$2^{4}$ $2^{6}$ 0.03 0.01 0.05 0.12 0.05 0.25
$2^{5}$ $2^{4}$ 0.05 0.02 0.05 0.14 0.06 0.25
$2^{5}$ $2^{5}$ 0.05 0.02 0.05 0.12 0.05 0.25
$2^{5}$ $2^{6}$ 0.05 0.02 0.05 0.13 0.04 0.25
$2^{6}$ $2^{4}$ 0.07 0.02 0.05 0.13 0.06 0.25
$2^{6}$ $2^{5}$ 0.06 0.02 0.05 0.14 0.08 0.26
$2^{6}$ $2^{6}$ 0.06 0.01 0.05 0.15 0.07 0.25
$2^{7}$ $2^{4}$ 0.05 0.02 0.05 0.17 0.07 0.25
$2^{7}$ $2^{5}$ 0.06 0.02 0.05 0.17 0.07 0.26
$2^{7}$ $2^{6}$ 0.07 0.02 0.05 0.17 0.08 0.24
$2^{8}$ $2^{4}$ 0.06 0.02 0.05 0.14 0.07 0.25
$2^{8}$ $2^{5}$ 0.06 0.02 0.05 0.15 0.05 0.24
$2^{8}$ $2^{6}$ 0.06 0.02 0.05 0.14 0.05 0.25
@@
Table 4: 1D case - 2-norm relative error between the exact source function $f(x)$ and its reconstruction, for different choices of the regularization parameter $\lambda$ and noise level $\varepsilon$. The fractional orders are $\beta=0.5$, $\omega=1.5$.
$\varepsilon=0.001$ $\varepsilon=0.01$
$\lambda$ $\lambda$
$n$ $S$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-4}$ $10^{-3}$ $10^{-2}$
$2^{4}$ $2^{4}$ 0.01 0.01 0.04 0.07 0.05 0.22
$2^{4}$ $2^{5}$ 0.01 0.01 0.04 0.08 0.06 0.23
$2^{4}$ $2^{6}$ 0.01 0.01 0.04 0.11 0.04 0.22
$2^{5}$ $2^{4}$ 0.04 0.01 0.04 0.15 0.07 0.23
$2^{5}$ $2^{5}$ 0.02 0.02 0.04 0.15 0.07 0.21
$2^{5}$ $2^{6}$ 0.04 0.02 0.04 0.14 0.05 0.23
$2^{6}$ $2^{4}$ 0.07 0.02 0.04 0.19 0.06 0.22
$2^{6}$ $2^{5}$ 0.07 0.02 0.04 0.19 0.04 0.22
$2^{6}$ $2^{6}$ 0.07 0.02 0.04 0.16 0.05 0.23
$2^{7}$ $2^{4}$ 0.08 0.02 0.04 0.23 0.07 0.22
$2^{7}$ $2^{5}$ 0.09 0.02 0.04 0.21 0.06 0.21
$2^{7}$ $2^{6}$ 0.08 0.02 0.04 0.19 0.06 0.22
$2^{8}$ $2^{4}$ 0.09 0.02 0.04 0.22 0.05 0.22
$2^{8}$ $2^{5}$ 0.09 0.03 0.04 0.19 0.05 0.23
$2^{8}$ $2^{6}$ 0.10 0.02 0.04 0.20 0.07 0.22
@@
Table 5: 1D case - 2-norm relative error between the exact source function $f(x)$ and its reconstruction, for different choices of the regularization parameter $\lambda$ and noise level $\varepsilon$. The fractional orders are $\beta=0.9$, $\omega=1.1$.
$\varepsilon=0.001$ $\varepsilon=0.01$
$\lambda$ $\lambda$
$n$ $S$ $10^{-5}$ $10^{-4}$ $10^{-3}$ $10^{-4}$ $10^{-3}$ $10^{-2}$
$2^{4}$ $2^{4}$ 0.01 0.01 0.04 0.06 0.05 0.23
$2^{4}$ $2^{5}$ 0.00 0.01 0.04 0.04 0.06 0.24
$2^{4}$ $2^{6}$ 0.01 0.00 0.04 0.04 0.05 0.24
$2^{5}$ $2^{4}$ 0.01 0.01 0.04 0.10 0.06 0.23
$2^{5}$ $2^{5}$ 0.01 0.01 0.04 0.10 0.05 0.23
$2^{5}$ $2^{6}$ 0.01 0.01 0.04 0.08 0.06 0.23
$2^{6}$ $2^{4}$ 0.03 0.02 0.04 0.16 0.06 0.24
$2^{6}$ $2^{5}$ 0.03 0.02 0.04 0.15 0.08 0.23
$2^{6}$ $2^{6}$ 0.03 0.02 0.04 0.17 0.06 0.23
$2^{7}$ $2^{4}$ 0.07 0.02 0.04 0.25 0.08 0.22
$2^{7}$ $2^{5}$ 0.06 0.03 0.04 0.23 0.09 0.23
$2^{7}$ $2^{6}$ 0.07 0.02 0.04 0.26 0.08 0.24
$2^{8}$ $2^{4}$ 0.10 0.03 0.04 0.31 0.09 0.22
$2^{8}$ $2^{5}$ 0.11 0.03 0.04 0.30 0.07 0.22
$2^{8}$ $2^{6}$ 0.11 0.03 0.04 0.28 0.07 0.23
6.2 Two dimensional setting
Now we proceed to the two-dimensional version of (4), setting
$\Omega=(0,\pi)^{2},\qquad T=1,$
$\gamma_{1}(x_{1},x_{2},t)=te^{x_{1}+x_{2}},\qquad\gamma_{2}(x_{1},x_{2},t)=t^{2}x_{1}x_{2},$
$q(t)=t^{2},\qquad\rho(x_{1},x_{2})=0.$
We construct the final time data $\mu(x_{1},x_{2})=v(x_{1},x_{2},T)$ by fixing
$f(x_{1},x_{2})=x_{1}x_{2}\sin(x_{1}x_{2})$
and numerically solving the corresponding direct problem, as we have done in Subsection 6.1. Likewise, we simulate the noisy data by adding a random quantity $\varepsilon\delta(x_{1},x_{2})$ to $\mu(x_{1},x_{2})$, with $\delta(x_{1},x_{2})$ randomly selected in $(-1,1)^{2}$ and $\varepsilon=0.01\cdot\|\mu\|_{2}$. Moreover, we always set $n_{1}=n_{2}=n$ for simplicity, so that ${\boldsymbol{n}}=(n,n)$.
6.2.1 Spectral analysis
We examine the effectiveness of $P_{{\boldsymbol{n}},S}$ in reducing the condition number and clustering the eigenvalues of the coefficient matrix at 1. The regularization parameter is set to $\lambda=5\cdot 10^{-3}$.
Figure 5 shows the eigenvalues in the complex plane for various choices of $\beta$ and $\omega$, with fixed problem dimensions $n=2^{4}$ (so that ${\boldsymbol{n}}=(2^{4},2^{4})$) and $S=2^{4}$. The plots are consistent with the theoretical discussion, showing that the eigenvalues accumulate near 0 for $A_{n,S}$. In the second and especially the third column, containing magnified graphs like in the one dimensional case, a tight cluster around 1 is clearly visible for the preconditioned matrix, with no eigenvalues near zero across all parameter settings.
Figure 5: 2D case - Eigenvalues of $A_{{\boldsymbol{n}},S}$ (left column) and $P_{{\boldsymbol{n}},S}^{-1}A_{{\boldsymbol{n}},S}$ (central column) in the complex plane, for different choices of the fractional orders, with ${\boldsymbol{n}}=(2^{4},2^{4})$ and $S=2^{4}$. For clarity, we also display magnified versions of the central graphs, focused on the cluster at 1 (right column).
Table 6 contains the 2-norm condition numbers of the nonpreconditioned and preconditioned coefficient matrix, for different choices of $\boldsymbol{n}$} and $S$. Once again, $A_{{\boldsymbol{n}},S}$ is extremely ill-conditioned, even for such small problem sizes. Similarly to the one dimensional case, $P_{{\boldsymbol{n}},S}$ is effective in reducing the condition number and performs very well when $A_{n,S}$ is particularly ill-conditioned, for $\beta\to 1$ and $\omega\to 2$.
Table 6: 2D case - 2-norm condition numbers of the nonpreconditioned and preconditioned coefficient matrix.
$\boldsymbol{n}$} $S$ - $P_{{\boldsymbol{n}},S}$ - $P_{{\boldsymbol{n}},S}$ - $P_{{\boldsymbol{n}},S}$
$2^{4}$ $2^{4}$ 34428 7 11123 65 18795 538
$2^{4}$ $2^{5}$ 69177 8 23296 98 54246 1153
$2^{5}$ $2^{4}$ 149171 9 27400 124 28649 1332
$2^{5}$ $2^{5}$ 315653 10 51158 187 80995 2922
6.2.2 GMRES method performance
We proceed to solving the linear system (11) with the GMRES method, reporting the results in Table 7. When the maximum number of iterations, which is the minimum between the size of the matrix and 1000, is reached before convergence, we simply write >1000>1000. The initial guess is set to the zero vector, the tolerance is $\texttt{tol}=10^{-8}$. The regularization parameter is set to $\lambda=5\cdot 10^{-3}$ and the noise level to $\varepsilon=0.01$.
In the columns related to the nonpreconditioned linear system, we observe consistently high iteration counts for all values of $\boldsymbol{n}$} and $S$, in accordance with the severe ill-conditioning. In fact, the number of iterations quickly becomes unmanageable, even using a coarse discretization.
In contrast, the columns associated with the preconditioned system show iteration counts that are consistent with the spectral analysis. $P_{{\boldsymbol{n}},S}$ delivers an excellent performance, particularly when $\beta$ approaches 1 and $\omega$ approaches 2, where the number of iterations remains very moderate and nearly independent from $\boldsymbol{n}$} and $S$. It is less effective when $\beta\to 0$ and $\omega\to 1$, in accordance with Table 6, which shows higher condition numbers for such values of the fractional orders.
Table 7: 2D case - Iterations required for solving the nonpreconditioned and preconditioned linear system with the GMRES method.
$\boldsymbol{n}$} $S$ - $P_{{\boldsymbol{n}},S}$ - $P_{{\boldsymbol{n}},S}$ - $P_{{\boldsymbol{n}},S}$
$2^{4}$ $2^{4}$ 867 10 460 20 486 41
$2^{4}$ $2^{5}$ >>1000 10 597 22 716 48
$2^{4}$ $2^{6}$ >>1000 10 795 25 >>1000 58
$2^{4}$ $2^{7}$ >>1000 11 >>1000 27 >>1000 68
$2^{4}$ $2^{8}$ >>1000 11 >>1000 30 >>1000 81
$2^{5}$ $2^{4}$ >>1000 11 861 25 760 60
$2^{5}$ $2^{5}$ >>1000 11 >>1000 29 >>1000 72
$2^{5}$ $2^{6}$ >>1000 12 >>1000 32 >>1000 87
$2^{5}$ $2^{7}$ >>1000 12 >>1000 35 >>1000 104
$2^{5}$ $2^{8}$ >>1000 13 >>1000 39 >>1000 124
$2^{6}$ $2^{4}$ >>1000 12 >>1000 32 >>1000 87
$2^{6}$ $2^{5}$ >>1000 13 >>1000 36 >>1000 105
$2^{6}$ $2^{6}$ >>1000 13 >>1000 40 >>1000 127